A monopoly firm faces a demand curve given by the following equation: P = $500 − 10Q, where Q equals quantity sold per day. Its marginal cost curve is MC = $100 per day. Assume that the firm faces no fixed cost. You may wish to arrive at the answers mathematically, or by using a graph (the graph is not required to be presented), either way, please provide a brief description of how you arrived at your result
b) How much will it charge?
c) Can you determine its profit per day? (Hint: you can; state how much it is.)
d) Suppose a tax of $1,000 per day is imposed on the firm. How will this affect its price?
e) How would the $1,000 per day tax its output per day?
f) How would the $1,000 per day tax affect its profit per day?
g) Now suppose a tax of $100 per unit is imposed. How will this affect the firm’s price?
h) How would a $100 per unit tax affect the firm’s profit maximizing output per day?
1
Expert's answer
2019-12-10T08:39:17-0500

b) MR = MC, so:

MR = TR'(Q) = 500 - 20Q, MC = 100,

500 - 20Q = 100,

Q = 20 units,

P = 500 - 10*20 = $300.

c) MC = ATC, so its profit per day is: TP = (P - MC)*Q = (300 - 100)*20 = $4,000.

d) If a tax of $1,000 per day is imposed on the firm, then its price will increase.

e) The $1,000 per day tax would decrease its output per day.

f) The $1,000 per day tax would decrease its profit per day.

g) If a tax of $100 per unit is imposed, then the firm’s price will increase.

h) A $100 per unit tax would decrease the firm’s profit maximizing output per day.